# A Hierarchical Framework for Measuring Scientific Paper Innovation via Large Language Models

### v1

@@ VI-D Comparison of Prompt Optimization Methods
VI-E LLM Computational Costs
Table XII compares the computational costs of question answering and score generation for LLaMA-3-8B, Mistral-7B, Qwen2.5-7B, DeepSeek-V3, and GPT-4o mini. The metrics include average output length (in characters) and average response time per input. For DeepSeek and GPT, the reported time only reflects API response time. All experiments were conducted on an NVIDIA RTX 3090 GPU (24 GB) without any inference-time optimization. We observe that large-scale models tend to generate more output tokens. Smaller LLMs (LLaMA-3-8B, Mistral-7B, Qwen2.5-7B) often suffer from repetition issues, leading to inflated output lengths. Among the open-source small-scale models, Mistral-7B shows competitive time efficiency. Combined with its strong innovation scoring performance in Section V-G, Mistral-7B stands out as a practical choice for HSPIM deployment.
TABLE XIII: Best question-prompt combinations found by HSPIMjoint
@@ VI-E LLM Computational Costs
VI-F Case Study
VI-F1 HSPIM Innovation Scores vs. Citations
VI-F2 Best Question-Prompt Combinations
Table XIII presents the best specific and common questions selected by DeepSeek-HSPIMjoint for each dataset. In terms of specific questions, we observe that certain question patterns appear consistently, especially for Abstract (“Is there clear evidence to support the contributions as novel?”) and Conclusion (“What specific innovative findings are emphasized in the research conclusions?”). These frequently selected question prompts suggest that innovation involves both identifying novelty and supporting the contributions with clear evidence. The specific questions are designed to align with the content of each section, enabling LLMs to extract relevant signals of contribution and novelty.
Additionally, the common question applies to all sections. It guides LLMs to assess both strengths and weaknesses of the paper. This helps reduce bias and ensures a balanced evaluation of innovation. From Table XIII, the common question “Provide arguments for both why this research may lack originality and why it could still be considered novel within its domain” appears twice. Overall, the genetic algorithm tends to select questions that ask authors to justify new ideas and reflect on their significance, which aligns with the goal of HSPIM to measure innovation with both novelty and contributions.
VII Discussion
@@
VII-A Addressing Challenges in Content-Based Innovation Measurement
@@
  • How to capture the full scope of innovation? Based on the innovation concept, HSPIM uses weighted novelty scoring to link section-level novelty to paper-level innovation. It uses confidence scores as weights to combine novelty scores. Table VIII shows the effectiveness of this weighted strategy. Section IV-B provides a theoretical explanation for the convergence of weighted scoring. In addition, QA augmentation helps to extract innovation-related information from text chunks for in-context learning. Table V and Table VI confirm its effectiveness in both quantifying innovation and improving interpretability.
  • How to improve model generalization? HSPIM builds on zero-shot LLM prompting. Table III shows that HSPIM outperforms both supervised deep learning baselines and LLM-based baselines. We use a genetic algorithm to optimize question prompts for improving robustness.
VII-B Limitations and Future Works
@@
VIII Conclusion
In this paper, we propose HSPIM, a hierarchical framework to measure the innovation of scientific papers. HSPIM is the first approach to conduct a Paper-to-Sections-to-QAs decomposition for assessing scientific innovation. HSPIM classifies text chunks by section type and transforms the complex task of measuring paper-level innovation into simpler sub-tasks of section-level question answering and weighted novelty score generation. HSPIM not only enables effective innovation quantification but also provides informative textual outputs, such as section-based QA pairs and scoring rationales, which are missing in previous content-based regression models. Extensive experiments demonstrate the effectiveness and interpretability of HSPIM in measuring innovation.

### v2

@@ 6.6 Impact of Section Misclassification
6.7 LLM Computational Costs
Table 17 compares the computational costs of question answering and score generation for LLaMA-3-8B, Mistral-7B, Qwen2.5-7B, DeepSeek-V3, and GPT-4o mini. The metrics include average output length (in characters) and average response time per input. For DeepSeek and GPT, the reported time only reflects API response time. All experiments were conducted on an NVIDIA RTX 3090 GPU (24 GB) without any inference-time optimization. We observe that large-scale models tend to generate more output tokens. Smaller LLMs (LLaMA-3-8B, Mistral-7B, Qwen2.5-7B) often suffer from repetition issues, leading to inflated output lengths. Among the open-source small-scale models, Mistral-7B shows competitive time efficiency. Combined with its strong innovation scoring performance in Section 5.7, Mistral-7B stands out as a practical choice for HSPIM deployment.
6.8 Case Study
6.8.1 HSPIM Innovation Scores vs. Citations
Table 18: Best question-prompt combinations found by HSPIMjoint.
@@ 6.8 Case Study
6.8.2 Best Question-Prompt Combinations
Table 18 presents the best specific and common questions selected by DeepSeek-HSPIMjoint for each dataset. In terms of specific questions, we observe that certain question patterns appear consistently, especially for Abstract (“Is there clear evidence to support the contributions as novel?”) and Conclusion (“What specific innovative findings are emphasized in the research conclusions?”). These frequently selected question prompts suggest that innovation involves both identifying novelty and supporting the contributions with clear evidence. The specific questions are designed to align with the content of each section, enabling LLMs to extract relevant signals of contribution and novelty.
Additionally, the common question applies to all sections. It guides LLMs to assess both strengths and weaknesses of the paper. This helps reduce bias and ensures a balanced evaluation of innovation. From Table 18, the common question “Provide arguments for both why this research may lack originality and why it could still be considered novel within its domain” appears twice. Overall, the genetic algorithm tends to select questions that ask authors to justify new ideas and reflect on their significance, which aligns with the goal of HSPIM to measure innovation with both novelty and contributions.
7 Discussion
@@
7.1 Addressing Challenges in Content-Based Innovation Measurement
  • 1.
@@ 7.1 Addressing Challenges in Content-Based Innovation Measurement
  • 2.
    How to capture the full scope of innovation? Based on the innovation concept, HSPIM uses weighted innovation scoring to link section-level novelty to paper-level innovation. It uses confidence scores as weights to measure each section’s contribution to overall innovation. Table 10 shows the effectiveness of this weighted strategy. Section 4.2 provides a theoretical explanation for the convergence of weighted scoring. In addition, QA augmentation helps to extract innovation-related information from text chunks for in-context learning. Table 5 and Table 8 confirm its effectiveness in both quantifying innovation and improving interpretability.
  • 3.
    How to improve model generalization? HSPIM builds on zero-shot LLM prompting. Table 3 shows that HSPIM outperforms both supervised deep learning baselines and LLM-based baselines. We use a genetic algorithm to optimize question prompts for improving robustness.
7.2 Limitations and Future Works
  • 1.
@@ 7.2 Limitations and Future Works
7.3 Practicality and Deployment
  • 1.
    Application potential: HSPIM provides a paper-level innovation score together with QA pairs and scoring reasons that can be inspected alongside the paper. The framework is training-free, applicable across domains, runs end to end, and works with both API-based and open source LLMs (see Section 3.2.5).
  • 2.
    Cost model: For a paper $p_{i}$ with $l_{i}$ chunks, the inference time is $\sum_{k=1}^{l_{i}}\tau_{\mathrm{QA},k}+\sum_{k=1}^{l_{i}}\tau_{\mathrm{score},k}$. Empirical times are reported in Table 17. Reducing the number of chunks lowers cost roughly in proportion to the reduction. In our experiments, HSPIMpruning could improve accuracy by using fewer section types. Besides, chunks are independent and can be processed in parallel to reduce total runtime. Prompt optimization is done offline and inference uses the selected prompts, keeping the framework training-free.
8 Conclusion
In this paper, we propose HSPIM, a hierarchical framework to measure the innovation of scientific papers. HSPIM is the first approach to conduct a Paper-to-Sections-to-QAs decomposition for assessing scientific innovation. HSPIM classifies text chunks by section type and transforms the complex task of measuring paper-level innovation into simpler sub-tasks of section-level question answering and weighted innovation scoring. Unlike prior content-based models that evaluate novelty from specific sections rather than the full paper, or peer review LLMs that generate review texts and attribute scores, HSPIM directly targets innovation measurement with a training-free hierarchical design. HSPIM not only enables effective innovation quantification but also provides informative textual outputs, including section-based QA pairs and scoring rationales, which are missing in previous content-based regression models. Extensive experiments demonstrate the effectiveness and interpretability of HSPIM in measuring innovation.